# Verifying Equilibria in Finite-Horizon Probabilistic Concurrent Game Systems

### v4

@@ 3.4. Strategy Improvements
The idea behind $\mathbb{G}_{i}\times\pi$ is to analyze the payoffs available to Agent $i$ when he is allowed to vary his strategy and deviate from $\pi_{i}$. Nevertheless, if we change Agent $i$’s output at every state $v\in V\times S\times[F]$, then we end up creating an exponentially large new strategy, placing us outside of PSPACE. The idea in this section is to prove the existence of a state $v^{*}$ such that, after only a “small, local” change to the output of Agent $i$ at $v^{*}$, the value of $v^{*}$ is immediately improved. Formally, this means that there is a relevant history reachable state $v^{*}\in R$ such that the strategy $\pi_{i}$ needs only to modify its output at $v^{*}$ to some deterministic distribution, i.e., a deterministic choice of action, in order to strictly increase the hitting probability $h_{v^{*}}$ of reaching $T$ from $v^{*}$. Connecting this back to $\mathbb{G}$, this means that only a very simple change needs to be made to $\pi_{i}$ in order to show that it is not optimal after the history $h$ corresponding to $v^{*}\in\mathbb{G}\times\pi$ is observed in $\mathbb{G}$.
@@ 3.5. Comparing the Two Probabilities
The arguments for $\mathbb{G}\times\pi$ and $\mathbb{G}\times\pi^{\prime}$ are the same, so WLOG, we work with $\mathbb{G}\times\pi$. Our proof uses backwards induction, so we start at the subset of states $\langle v,s,n|n=F\rangle$. For a state $x\in\langle v,s,n|n=F\rangle$, the hitting probability $h_{x}$ is either 00 or $1$ depending on whether it is a target state or not.
Now consider the set of states $\langle v,s,n|n=F-1\rangle$. For a state $x$ in $\langle v,s,n|n=F-1\rangle$ the hitting probability $h_{x}$ can be calculated by the equation $h_{x}=p_{a}\cdot h_{a}+p_{b}\cdot h_{b}\ldots$ where the hitting probabilities $h_{a}$ and $h_{b}$ on the right-hand side (RHS) refer to states in the set $\langle v,s,n|n=F\rangle$. Recall the definitions of $b$, the concurrency bound, $\mathbb{L}$, the size bound of probabilities, and $A$, the set of actions from Definition 2 and the definition of $\mathbb{G}\times\pi$ from Section 3. We know the following facts about the RHS :
@@
    The probabilities $p_{i}$ of transitioning from the state $\langle v_{1},s_{1},n_{1}\rangle$ to $\langle v_{2},s_{2},n_{2}\rangle$ are either 00 or given by the expression $\sum_{\theta\in supp(O(s_{1},v_{1}))}\mathbb{P}(O(s_{1},v_{1})[\theta])\cdot\delta(v_{1},\theta,v_{2})$, as shown at the start of Section 3. The term $\mathbb{P}(O(s_{1},v_{1})[\theta])$ refers to the probability that the product transducer $T_{\pi}$, as given in Definition 2, assigns to $\theta$ on an input of $v_{1}$ in state $s_{1}$. This, itself, is the product of at most $b$ probabilities from the individual transducers $\pi_{i}$. Each probability in each transducer $\pi_{i}$ is exactly $\mathbb{L}$ bits in length. Therefore, $\mathbb{P}(O(s_{1},v_{1})[\theta])$ is at most $b\cdot\mathbb{L}$ bits in length. Furthermore, we multiply this probability by $\delta(v_{1},\theta,v_{2})$, which is also $\mathbb{L}$ bits in length, giving us a total of $(b+1)\cdot\mathbb{L}$ bits. Finally, we have a sum over $supp(O(s_{1},v_{1}))$, which has an upper bound of $|A|^{b}$ elements. Thus, the total number of bits used to represent a transition probability $p_{i}$ in $\mathbb{G}\times\pi$ is $\log(|A|^{b})+(b+1)\cdot\mathbb{L}$. We refer to this number as $\mathbb{L}_{\mathbb{G}\times\pi}$, the size of the representation of the probabilities in $\mathbb{G}\times\pi$. As we have already shown in Lemma 1, $\mathbb{L}_{\mathbb{G}\times\pi}$ is polynomial in the size of the input.
Let us consider the equation $h_{x}=p_{a}\cdot h_{a}+p_{b}\cdot h_{b}\ldots$. We claim that the size of the representation of the evaluation of the RHS has an upper bound of $\log(|V|\cdot|S|)+\mathbb{L}_{\mathbb{G}\times\pi}$. Each term $h_{i}\cdot p_{i}$ has a representation of size $\mathbb{L}_{\mathbb{G}\times\pi}$, since $h_{i}$ is either 00 or $1$. Also, since there are at most $|V|\cdot|S|$ terms on the RHS, we are at most considering a sum over $|V|\cdot|S|$ elements, which can be upper bounded by multiplying by $|V|\cdot|S|$, adding at most $\log(|V|\cdot|S|)$ bits. Therefore, it is possible to represent the hitting probability $h_{x}$ for a state $x\in\langle v,s,n|n=F-1\rangle$ using at most $\log(|V|\cdot|S|)+\mathbb{L}_{\mathbb{G}\times\pi}$ bits.
@@ 3.6.2. Full Concurrency Model
In Section 3, we constructed a Markov Chain $\mathbb{G}\times\pi$ with the critical property that elements of its transition table $P$ could be queried on-the-fly in PSPACE. We now show that the elements of $P$ – representing the probability that a state $\langle v_{1},s_{1},n_{1}\rangle$ transitions to a state $\langle v_{2},s_{2},n_{2}\rangle$ – could be computed in PSPACE even when a full concurrency model is considered as opposed to the $b$-bounded one presented in this paper. Recall that elements of $P$ were given by a product $p_{v}\cdot p_{s}\cdot p_{n}$. In a full concurrency model, $p_{s}$ would not change since it considered every strategy $\pi_{i}$ to begin with. Additionally, $p_{n}$ would not change as well since it only accounted for the time index. The only change would be to $p_{v}$, which was calculated through the sum $\sum_{\theta\in supp(O(s_{1},v_{1}))}\mathbb{P}(O(s_{1},v_{1})[\theta])\cdot\delta(v_{1},\theta,v_{2})$. This sum ranged over $supp(O(s_{1},v_{1}))$, which could only contain polynomially many elements w.r.t the input due to the $b$-bounded nature of the game. However, even if it ranged over $\Theta=\bigtimes_{i\in\Omega}A_{i}$, the cross-product of every action set as in a fully concurrent setting, this would still only be an exponential size set. It could, therefore, be iterated over in PSPACE, meaning that $p_{v}$ could still be computed in PSPACE. The rest of the results would then follow in exactly the same manner, giving us an analogous statement of Theorem 7 for the subgame perfect equilibrium verification problem in full concurrency models. Similarly, the EXPTIME upper bound presented in Theorem 8 for the Nash equilibrium verification problem would also be preserved in full concurrency models.
@@ 3.6.3. Probabilistic Transition Functions
We now consider what would happen if we were to consider a transducer model with a probabilistic transition function as opposed to a deterministic one. In $\mathbb{G}\times\pi$ as defined in Section 3, the probability that a state $\langle v_{1},s_{1},n_{1}\rangle$ transitions to a state $\langle v_{2},s_{2},n_{2}\rangle$ could still be computed in PSPACE. Recall that this probability was given as the product of three terms, $p_{v}\cdot p_{s}\cdot p_{n}$. The only change to this would be to $p_{s}$, the transition probability between $s_{1}=\langle s_{1}^{1}\ldots s_{1}^{k}\rangle$ and $s_{2}=\langle s_{2}^{1}\ldots s_{2}^{k}\rangle$. This probability would then be given by the product ×$\bigtimes_{i\in\Omega}\mathbb{P}(\gamma^{i}(s^{i}_{1},v_{1})[s^{i}_{2}])$, where $\mathbb{P}(\gamma^{i}(s^{i}_{1},v_{1})[s^{i}_{2}])$ is the probability the distribution $\gamma^{i}(s^{i}_{1},v_{1})$ assigns to $s^{i}_{2}$. This product could be computed in PSPACE as well; if $m$ is the size of the largest encoding of a probability in $\pi$, then the encoding of the product would be at most $\log(2^{m\cdot k})$ bits, which is polynomial w.r.t the input.
@@ 4.1. Upper Bound
Broadly speaking, an EXPTIME upper bound for the verification problem appears in various other works in various other forms as a “folk theorem,” e.g. [YSD+22]. In this work, we make the details explicit and consider issues of numerical issues that are often ignored in other works.
@@ 4.2. Lower Bound
[…]
@@
    $\kappa:R\rightarrow\{\mathrm{accept},\mathrm{reject},\vee,\wedge,-\}$ is the state labeling function that associates each state with its type. A state can be accepting, rejecting, existential ($\vee$), universal ($\wedge$), or deterministic (--).
@@
Alternating Turing machines generalize non-deterministic Turing machines by classifying machine states as existential ($\vee$), universal ($\wedge$), or deterministic (--). Existential states work exactly as described previously in the description of non-deterministic Turing machines. In an existential state, multiple transitions are possible, and the machine can choose the one that best serves the purpose of acceptance. The dual of this concept is the universal state, in which multiple transitions are possible, but instead of choosing one, the computation must be accepting from all successor states. Deterministic transitions work as one would expect from a deterministic Turing machine. While it is possible to encapsulate deterministic behavior by creating either an alternating or universal state with a single successor, it is more convenient notationally to introduce a special category for deterministic states.
@@ 4.2.1. A High Level View of the Reduction
This problem is hard for the class APSPACE = EXPTIME [CS76]. In this decision problem, the machine $M$ can never use more than $n$ cells in an accepting computation, as any computation that tries to access more than that amount of space is immediately rejected. Since the machine can never use more than $n$ cells, we introduce the concept of an ID – a string that represents the complete information of the tape and machine at a given moment in time. {exa}For example, an ID may read $122\langle r_{3},5\rangle 68$. This string conveys the following information : (1) The string is $6$ characters long – $\langle r_{3},5\rangle$ is considered to be a single character. This means that the bounded Turing machine can only use $6$ cells, i.e., $n=6$. (2) The first cell, counting from the left, reads `$1^{\prime}$. For this example, $\Gamma$ is given by the set of digits. (3) The head of the Turing machine is currently reading the fourth cell from the left, represented by the character $\langle r_{3},5\rangle$. The machine is in state $r_{3}\in R$, reading the character ${}^{\prime}5^{\prime}$ written on the fourth cell from the left.
@@
In order to update the $i$-the character in the current ID, Agent $i$ needs global information about the transitions in $M$. This information is stored in the states in $V$. Intuitively, when the head of $M$ is positioned over cell $i$, the strategy $\pi_{i}$ outputs an action that describes the complete information of the transition – which character was written, which cell the head was just at (in this case, cell $i$), which state the machine has transitioned to, and which direction the head is moving in. Once Agent $i$ outputs this action, the game $\mathbb{G}$ deterministically transitions to a state that is labeled by this same information, allowing all agents to view and update the internal state of their strategy representing their ID character. For example, if Agent $i-1$ sees that the head was previously at cell $i$ and is moving left in state $q$, they know the head will be positioned over their cell in state $q$ next and can therefore update the state of $\pi_{i-1}$ accordingly. The state of $V$ then deterministically transitions to a state indicating that the head is over cell $i-1$ and it is awaiting an action from Agent $i-1$.
While this scheme works for deterministic transitions, the alternating Turing machine $M$ has universal and existential transitions as well. The high-level idea behind universal transitions is to output each transition with .5.5.5.5 probability. Therefore, in order for the simulation of the machine to reach an accepting ID configuration with probability $1$, it must be the case that $M$ has an accepting computation tree, e.g., a computation tree where every branch is accepting. The high-level idea behind existential transitions is that Agent $n+1$ is allowed to choose the existential transitions. If the machine $M$ is an existential state over the cell $i$, it allows Agent $n+1$ to choose which of the two possibles successor should be taken. This information is once again encoded into the state space $V$, and Agent $i$ uses this new information when $\pi_{i}$ outputs their transition action.
@@
Before moving on to the technical details, we provide a brief overview of the reduction idea. The strategy $\pi$ tells Agent $n+1$ to take the fixed payoff at the start of the game. Therefore, the machine $M$ is not being simulated, and Agent $n+1$ is not deciding between the $\alpha$ and $\beta$ successors in the simulation of $M$. This fixed payoff is chosen to be higher than the payoff obtained by any non-accepting $M$, so if $M$ does not accept, then $\pi$ is a Nash equilibrium. If, however, the machine $M$ has an accepting computation tree, then there is a strategy for Agent $n+1$ to simulate the machine $M$ and receive a payoff of $1$, meaning that $\pi$ is not a Nash equilibrium.

### v5

@@ 3.4. Strategy Improvements
The idea behind $\mathbb{G}_{i}\times\pi$ is to analyze the payoffs available to Agent $i$ when he is allowed to vary his strategy and deviate from $\pi_{i}$. Nevertheless, if we change Agent $i$’s output at every state $v\in V\times S\times[F]$, then we end up creating an exponentially large new strategy, placing us outside of PSPACE. The idea in this section is to prove the existence of a state $v^{*}$ such that, after only a “small, local” change to the output of Agent $i$ at $v^{*}$, the value of $v^{*}$ is immediately improved. Formally, this means that there is a relevant history-reachable state $v^{*}\in R$ such that the strategy $\pi_{i}$ needs only to modify its output at $v^{*}$ to some deterministic distribution, i.e., a deterministic choice of action, in order to strictly increase the hitting probability $h_{v^{*}}$ of reaching $T$ from $v^{*}$. Connecting this back to $\mathbb{G}$, this means that only a very simple change needs to be made to $\pi_{i}$ in order to show that it is not optimal after the history $h$ corresponding to $v^{*}\in\mathbb{G}\times\pi$ is observed in $\mathbb{G}$.
@@ 3.5. Comparing the Two Probabilities
The arguments for $\mathbb{G}\times\pi$ and $\mathbb{G}\times\pi^{\prime}$ are the same, so WLOG, we work with $\mathbb{G}\times\pi$. Our proof uses backwards induction, so we start at the subset of states $\langle v,s,n|n=F\rangle$. For a state $x\in\langle v,s,n|n=F\rangle$, the hitting probability $h_{x}$ is either 0 or $1$ depending on whether it is a target state or not. The intuitive idea behind this proof is to define subsets $\langle v,s,n|n=k\rangle$ for $k\in[F]$ and show that when moving backwards through the index $n$, the length of the representation for the hitting probabilities grows additively. Since there are exponentially many such subsets (since $|F|$ is exponential), we end up adding exponentially many times, giving us an exponential-sized result.
Now, consider the set of states $\langle v,s,n|n=F-1\rangle$. For a state $x$ in $\langle v,s,n|n=F-1\rangle$ the hitting probability $h_{x}$ can be calculated by the equation $h_{x}=p_{a}\cdot h_{a}+p_{b}\cdot h_{b}\ldots$ where the hitting probabilities $h_{a}$ and $h_{b}$ on the right-hand side (RHS) refer to states in the set $\langle v,s,n|n=F\rangle$. Recall the definitions of $b$, the concurrency bound, $\mathbb{L}$, the size bound of probabilities, and $A$, the set of actions from Definition 2 and the definition of $\mathbb{G}\times\pi$ from Section 3. We know the following facts about the RHS :
@@
    The probabilities $p_{i}$ of transitioning from the state $\langle v_{1},s_{1},n_{1}\rangle$ to $\langle v_{2},s_{2},n_{2}\rangle$ are either 0 or given by the expression $\sum_{\theta\in supp(O(s_{1},v_{1}))}\mathbb{P}(O(s_{1},v_{1})[\theta])\cdot\delta(v_{1},\theta,v_{2})$, as shown at the start of Section 3. The term $\mathbb{P}(O(s_{1},v_{1})[\theta])$ refers to the probability that the product transducer $T_{\pi}$, as given in Definition 2, assigns to $\theta$ on an input of $v_{1}$ in state $s_{1}$. This, itself, is the product of at most $b$ probabilities from the individual transducers $\pi_{i}$. Each probability in each transducer $\pi_{i}$ is exactly $\mathbb{L}$ bits in length. Therefore, $\mathbb{P}(O(s_{1},v_{1})[\theta])$ is at most $b\cdot\mathbb{L}$ bits in length. Furthermore, we multiply this probability by $\delta(v_{1},\theta,v_{2})$, which is also $\mathbb{L}$ bits in length, giving us a total of $(b+1)\cdot\mathbb{L}$ bits. Finally, we have a sum over $supp(O(s_{1},v_{1}))$, which has an upper bound of $|A|^{b}$ elements. Thus, the total number of bits used to represent a transition probability $p_{i}$ in $\mathbb{G}\times\pi$ is $\log(|A|^{b})+(b+1)\cdot\mathbb{L}$. We refer to this number as $\mathbb{L}_{\mathbb{G}\times\pi}$, the size of the representation of the probabilities in $\mathbb{G}\times\pi$. As we have already shown in Lemma 1, $\mathbb{L}_{\mathbb{G}\times\pi}$ is polynomial in the size of the input.
Let us consider the equation $h_{x}=p_{a}\cdot h_{a}+p_{b}\cdot h_{b}\ldots$. We claim that the size of the representation of the evaluation of the RHS has an upper bound of $\log(|V|\cdot|S|)+\mathbb{L}_{\mathbb{G}\times\pi}$. Each term $h_{i}\cdot p_{i}$ has a representation of size $\mathbb{L}_{\mathbb{G}\times\pi}$, since $h_{i}$ is either 0 or $1$. Also, since there are at most $|V|\cdot|S|$ terms on the RHS, we are at most considering a sum over $|V|\cdot|S|$ elements, which can be upper bounded by multiplying by $|V|\cdot|S|$, adding at most $\log(|V|\cdot|S|)$ bits. Therefore, it is possible to represent the hitting probability $h_{x}$ for a state $x\in\langle v,s,n|n=F-1\rangle$ using at most $\log(|V|\cdot|S|)+\mathbb{L}_{\mathbb{G}\times\pi}$ bits.
@@ 3.6.2. Full Concurrency Model
In Section 3, we constructed a Markov Chain $\mathbb{G}\times\pi$ with the critical property that elements of its transition table $P$ could be queried on-the-fly in PSPACE. We now show that the elements of $P$ – representing the probability that a state $\langle v_{1},s_{1},n_{1}\rangle$ transitions to a state $\langle v_{2},s_{2},n_{2}\rangle$ – could be computed in PSPACE even when a full concurrency model is considered as opposed to the $b$-bounded one presented in this paper. Recall that elements of $P$ were given by a product $p_{v}\cdot p_{s}\cdot p_{n}$. In a full concurrency model, $p_{s}$ would not change since it considers every strategy $\pi_{i}$ to begin with. Additionally, $p_{n}$ would not change as well since it only accounted for the time index. The only change would be to $p_{v}$, which was calculated through the sum $\sum_{\theta\in supp(O(s_{1},v_{1}))}\mathbb{P}(O(s_{1},v_{1})[\theta])\cdot\delta(v_{1},\theta,v_{2})$. This sum ranged over $supp(O(s_{1},v_{1}))$, which could only contain polynomially many elements w.r.t the input due to the $b$-bounded nature of the game. Even if, however, it ranged over $\Theta=\bigtimes_{i\in\Omega}A_{i}$, the cross-product of every action set as in a fully concurrent setting, this would still only be an exponential-size set. It could, therefore, be iterated over in PSPACE, meaning that $p_{v}$ could still be computed in PSPACE. The rest of the results would then follow in exactly the same manner, giving us an analogous statement of Theorem 7 for the subgame perfect equilibrium verification problem in full concurrency models. Similarly, the EXPTIME upper bound presented in Theorem 8 for the Nash equilibrium verification problem would also be preserved in full concurrency models.
@@ 3.6.3. Probabilistic Transition Functions
We now consider what would happen if we were to consider a transducer model with a probabilistic transition function as opposed to a deterministic one. In $\mathbb{G}\times\pi$ as defined in Section 3, the probability that a state $\langle v_{1},s_{1},n_{1}\rangle$ transitions to a state $\langle v_{2},s_{2},n_{2}\rangle$ could still be computed in PSPACE. Recall that this probability was given as the product of three terms, $p_{v}\cdot p_{s}\cdot p_{n}$. The only change to this would be to $p_{s}$, the transition probability between $s_{1}=\langle s_{1}^{1}\ldots s_{1}^{k}\rangle$ and $s_{2}=\langle s_{2}^{1}\ldots s_{2}^{k}\rangle$. This probability would then be given by the product $\bigtimes_{i\in\Omega}\mathbb{P}(\gamma^{i}(s^{i}_{1},v_{1})[s^{i}_{2}])$, where $\mathbb{P}(\gamma^{i}(s^{i}_{1},v_{1})[s^{i}_{2}])$ is the probability the distribution $\gamma^{i}(s^{i}_{1},v_{1})$ assigns to $s^{i}_{2}$. This product could be computed in PSPACE as well; if $m$ is the size of the largest encoding of a probability in $\pi$, then the encoding of the product would be at most $\log(2^{m\cdot k})$ bits, which is polynomial w.r.t the input.
@@ 4.1. Upper Bound
Broadly speaking, an EXPTIME upper bound for the verification problem appears in various other works in various other forms as a “folk theorem,” e.g. [YSD+22]. In this work, we make the details explicit and consider issues of numerical representation that are often ignored in other works.
@@ 4.2. Lower Bound
    $\Delta:R\times\Gamma\rightarrow\mathcal{P}(R\times\Gamma\times\{\rightarrow,\leftarrow\})$ is the transition function, where $\mathcal{P}(\cdot)$ refers to the powerset construction. The set $\{\rightarrow,\leftarrow\}$ represents the left and right directions – how the head moves over the tape in the Turing machine. Since this is an alternating Turing machine, the range of the transition function is a powerset representing all possible transitions available from a single state-character pair. For notational convenience, we assume that the range of $\Delta$ on an arbitrary input has at most size two, i.e., there are only two possible successors from each state-character pair. This assumption does not change the complexity-theoretic properties of the decision problem from which we wish to reduce.
@@
    $\kappa:R\rightarrow\{\mathrm{accept},\mathrm{reject},\vee,\wedge,-\}$ is the state labeling function that associates each state with its type. A state can be accepting, rejecting, existential ($\vee$), universal ($\wedge$), or deterministic (-).
@@
Alternating Turing machines generalize non-deterministic Turing machines by classifying machine states as existential ($\vee$), universal ($\wedge$), or deterministic (-). Existential states work exactly as described previously in the description of non-deterministic Turing machines. In an existential state, multiple transitions are possible, and the machine can choose the one that best serves the purpose of acceptance. The dual of this concept is the universal state, in which multiple transitions are possible, but instead of choosing one, the computation must be accepting from all successor states. Deterministic transitions work as one would expect from a deterministic Turing machine. While it is possible to encapsulate deterministic behavior by creating either an alternating or universal state with a single successor, it is more convenient notationally to introduce a special category for deterministic states.
@@ 4.2.1. A High Level View of the Reduction
This problem is hard for the class APSPACE = EXPTIME [CS76]. In this decision problem, the machine $M$ can never use more than $n$ cells in an accepting computation, as any computation that tries to access more than that amount of space is immediately rejected. Since the machine can never use more than $n$ cells, we introduce the concept of an ID – a string that represents the complete information of the tape and machine at a given moment in time. {exa}For example, an ID may read $122\langle r_{3},5\rangle 68$. This string conveys the following information : (1) The string is $6$ characters long – $\langle r_{3},5\rangle$ is considered to be a single character. This means that the bounded Turing machine can only use $6$ cells, i.e., $n=6$. (2) The first cell, counting from the left, reads $1^{\prime}$. For this example, $\Gamma$ is given by the set of digits. (3) The head of the Turing machine is currently reading the fourth cell from the left, represented by the character $\langle r_{3},5\rangle$. The machine is in state $r_{3}\in R$, reading the character $^{\prime}5^{\prime}$ written on the fourth cell from the left.
@@
In order to update the $i$-the character in the current ID, Agent $i$ needs global information about the transitions in $M$. This information is stored in the states in $V$. Intuitively, when the head of $M$ is positioned over cell $i$, the strategy $\pi_{i}$ outputs an action that describes the complete information of the transition – which character was written, which cell the head was just at (in this case, cell $i$), which state the machine has transitioned to, and which direction the head is moving in. Once Agent $i$ outputs this action, the game $\mathbb{G}$ deterministically transitions to a state that is labeled by this same information, allowing all agents to view and update the internal state of their strategy representing their ID character. For example, if Agent i1i-1 sees that the head was previously at cell $i$ and is moving left in state $q$, they know the head will be positioned over their cell in state $q$ next and can therefore update the state of $\pi_{i-1}$ accordingly. The state of $V$ then deterministically transitions to a state indicating that the head is over cell i1i-1 and it is awaiting an action from Agent i1i-1.
While this scheme works for deterministic transitions, the alternating Turing machine $M$ has universal and existential transitions as well. The high-level idea behind universal transitions is to output each transition with .5.5 probability. Therefore, in order for the simulation of the machine to reach an accepting ID configuration with probability $1$, it must be the case that $M$ has an accepting computation tree, e.g., a computation tree where every branch is accepting. The high-level idea behind existential transitions is that Agent $n+1$ is allowed to choose the existential transitions. If the machine $M$ is an existential state over the cell $i$, it allows Agent $n+1$ to choose which of the two possibles successor should be taken. This information is once again encoded into the state space $V$, and Agent $i$ uses this new information when $\pi_{i}$ outputs their transition action.
@@
Before moving on to the technical details, we provide a brief overview of the reduction idea. The strategy $\pi$ tells Agent $n+1$ to take the fixed payoff at the start of the game. Therefore, the machine $M$ is not being simulated, and Agent $n+1$ is not deciding between the $\alpha$ and $\beta$ successors in the simulation of $M$. This fixed payoff is chosen to be higher than the payoff obtained by any non-accepting $M$, so if $M$ does not accept, then $\pi$ is a Nash equilibrium. If, however, the machine $M$ has an accepting computation tree, then there is a strategy for Agent $n+1$ to simulate the machine $M$ and receive a payoff of $1$, meaning that $\pi$ is not a Nash equilibrium. This is the central idea behind the reduction, and it is illustrated in Figure 3.
[…]
Figure 3. The crucial mechanism of the reduction, Agent $n+1$’s first choice. The Agent can either choose the fixed payoff mechanism with $\alpha$, or start the probabilistic simulation of $M$ by the other $n$ agents with $\beta$ with the intention of reaching an accepting configuration. Since $M$ is an alternating Turing machine, it admits a computation tree. If Agent $n+1$ chooses $\beta$, the game proceeds by probabilistically sampling one of the branches in the computation tree. The numbers are chosen in such a way that if even one branch in the computation tree is nonaccepting (meaning $M$ is nonaccepting), Agent $n+1$ gets a better chance of reaching their goal through choosing $\alpha$ and taking the fixed payoff mechanism.